# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 3. Slice Theorem for cohomogeneity one geodesic spaces
    $\mathcal{T}_{u}^{\ast}=\mathcal{T}_{u^{\ast}}$,
@@
    $\Gamma_{u^{\ast}}(x^{\ast})=\Gamma_{u}(x)^{\ast}$ and $\Gamma^{-}_{u^{\ast}}(x^{\ast})=\Gamma^{-}_{u}(x)^{\ast}$ for every $x\in X$,
@@
    $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$ and $A_{-}(u)^{\ast}=A_{-}(u^{\ast})$.
@@
    $(\mathcal{T}^{nb}_{u})^{\ast}=\mathcal{T}^{nb}_{u^{\ast}}$.
Moreover, if $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$, then $\mathcal{T}^{nb}_{u^{\ast}}$ is equal to $X^{\ast}$ in the first three cases or $\mathbf{S}^{1}\setminus\left\{-x_{0}^{\ast}\right\}$ respectively, where $-x_{0}^{\ast}$ denotes the antipodal orbit to $x_{0}^{\ast}$.
@@
First we prove that $\mathcal{T}_{u}^{\ast}\subseteq\mathcal{T}_{u^{\ast}}$. Consider $x^{\ast}\in\mathcal{T}^{\ast}_{u}$, for some $x\in\mathcal{T}_{u}$. Then there exists $y\in X$ with $x\neq y$ and $(x,y)\in\Gamma_{u}\cup\Gamma^{-}_{u}$. We then have two cases depending on whether $(x,y)\in\Gamma_{u}$ or $(x,y)\in\Gamma^{-}_{u}$.
Assume first that $(x,y)\in\Gamma_{u}$. Then, by Lemma 3.2, $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Now, suppose, by contradiction, that $x^{\ast}=y^{\ast}$, that is, $y=gx$ for some $g\in G$. Then we have that
$d(x,y)=d(x,gx)=d(G(x_{0}),x)-d(G(x_{0}),gx)=0.$
Thus $y=x$, which is a contradiction. Hence $y\not\in G(x)$, and therefore $x^{\ast}\neq y^{\ast}$. Thus we have that $x^{\ast}\in\mathcal{T}_{u^{\ast}}$. The proof in the case that $(x,y)\in\Gamma^{-}_{u}$ is analogous. Thus, we have shown that $\mathcal{T}_{u}^{\ast}\subseteq\mathcal{T}_{u^{\ast}}$.
To prove the opposite inclusion, now consider $x^{\ast}\in\mathcal{T}_{u^{\ast}}$. Then there is $y^{\ast}\in X^{\ast}$ such that $y^{\ast}\neq x^{\ast}$ and $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}\cup\Gamma_{u^{\ast}}^{-}$. Then by Lemma 3.2, for some representatives $\tilde{x}\in G(x)$ of $x^{\ast}$ and $\tilde{y}\in G(y)$ of $y^{\ast}$, we have $(\tilde{x},\tilde{y})\in R_{u}$: We assume first that $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$. Then it follows that for some $g\in G$ we have $(gx,y)\in\Gamma_{u}$. In the case where $(x^{\ast},y^{\ast})\in\Gamma^{-}_{u^{\ast}}$ we have then that there exists $g^{\prime}\in G$ such that $(x,g^{\prime}y)\in\Gamma_{u}^{-}$. Observe that if $\tilde{x}=\tilde{y}$ we have that $G(x)\cap G(y)\neq\varnothing$, which is a contradiction. Thus we conclude that $\tilde{x}\in\mathcal{T}_{u}$, whence $\mathcal{T}_{u}^{\ast}\supseteq\mathcal{T}_{u^{\ast}}$. Thus we have that $\mathcal{T}_{u}^{\ast}=\mathcal{T}_{u^{\ast}}$, as claimed.
We now show that $\Gamma_{u^{\ast}}(x^{\ast})\subseteq\Gamma_{u}(x)^{\ast}$. Consider $y^{\ast}\in\Gamma_{u^{\ast}}(x^{\ast})$. This implies that there exist representatives $\tilde{x}\in G(x)$ of $x^{\ast}$ and $\tilde{y}\in G(y)$ of $y^{\ast}$ such that $(\tilde{x},\tilde{y})\in\Gamma_{u}$. In particular $\tilde{y}\in\Gamma_{u}(\tilde{x})$. Now assume that $\tilde{x}=gx$ for some $g\in G$, then $g^{-1}\tilde{y}\in\Gamma_{u}(x)$. Thus we conclude that $y^{\ast}\in\Gamma_{u}(x)^{\ast}$. This shows that $\Gamma_{u^{\ast}}(x^{\ast})\subseteq\Gamma_{u}(x)^{\ast}$.
For the other inclusion, given $y\in\Gamma_{u}(x)$, we have that $(x,y)\in\Gamma_{u}$. Thus $(x^{\ast},y^{\ast})\in\Gamma_{u^{\ast}}$ by Lemma 3.2 and we conclude that $y^{\ast}\in\Gamma_{u^{\ast}}(x^{\ast})$. Thus the claim that $\Gamma_{u^{\ast}}(x^{\ast})=\Gamma_{u}(x)^{\ast}$ for every $x\in X$ is proved. An analogous argument shows that $\Gamma^{-}_{u^{\ast}}(x^{\ast})=\Gamma^{-}_{u}(x)^{\ast}$ for every $x\in X$.
Now we show that $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$ and $A_{-}(u)^{\ast}=A_{-}(u^{\ast})$. We prove this for $A_{+}(u)$ since the argument for $A_{-}(u)$ is analogous.
Consider $y^{\ast}\in A_{+}(u^{\ast})$. Then there exists $z^{\ast},w^{\ast}\in\Gamma_{u^{\ast}}(y^{\ast})$ such that $(z^{\ast},w^{\ast})\not\in R_{u^{\ast}}$. Fix a representative $\tilde{y}\in G(y)$ of $y^{\ast}$. Then there exist representatives $\tilde{z}\in G(z)$ of $z^{\ast}$ and $\tilde{w}\in G(w)$ of $w^{\ast}$ with $\tilde{z},\tilde{w}\in\Gamma_{u}(\tilde{y})$. If we assume that $(\tilde{z},\tilde{w})\in R_{u}$ it would follow from Lemma 3.2 that $(z^{\ast},w^{\ast})\in R_{u^{\ast}}$, which is a contradiction. Thus we have that $\tilde{z},\tilde{w}\in\Gamma_{u}(\tilde{y})$ and $(\tilde{z},\tilde{w})\not\in R_{u}$, i.e. $\tilde{y}\in A_{+}(u)$. This implies that $y^{\ast}\in A_{+}(u)^{\ast}$, whence, $A_{+}(u^{\ast})\subseteq A_{+}(u)^{\ast}$.
To prove the opposite inclusion, given $y^{\ast}\in A_{+}(u)^{\ast}$ for some $y\in A_{+}(u)$, we have that there exists $z,w\in\Gamma_{u}(y)$ with $(z,w)\not\in R_{u}$. Then $z^{\ast},w^{\ast}\in\Gamma_{u^{\ast}}(y^{\ast})$ and by Lemma 3.2 we obtain $(z^{\ast},w^{\ast})\not\in R_{u^{\ast}}$, i.e. $y^{\ast}\in A_{+}(u^{\ast})$. Thus we have shown that $A_{+}(u)^{\ast}\subseteq A_{+}(u^{\ast})$, and in turn that $A_{+}(u)^{\ast}=A_{+}(u^{\ast})$.
Now it readily follows from the definition of the non-branching transport set and the previous paragraphs that
$\mathcal{T}^{nb}_{u^{\ast}}=\mathcal{T}_{u^{\ast}}\setminus\left(A_{+}(u^{\ast})\cup A_{-}(u^{\ast})\right)=\mathcal{T}_{u}^{\ast}\setminus(A_{+}(u)^{\ast}\cup A_{-}(u)^{\ast})=(\mathcal{T}^{nb}_{u})^{\ast},$
obtaining the desired conclusion.
For the final claim of the Lemma we only explain the case when $X^{\ast}$ is isometric to $[-1,1]$ and $x_{0}^{\ast}=-1$ as all other cases are very similar. Let us observe that when $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$ we have that if $x_{0}^{\ast}=-1$, then $u^{\ast}(x^{\ast})-u^{\ast}(y^{\ast})=d^{\ast}(x^{\ast},y^{\ast})$ if and only if $y^{\ast}\leq x^{\ast}$. Thus we conclude that,
$\Gamma_{u^{\ast}}$ =$=\{(s,t)\in[-1,1]\times[-1,1]\mid t\leq s\},$
$\Gamma_{u^{\ast}}^{-}$ =$=\{(s,t)\in[-1,1]\times[-1,1]\mid s\leq t\},$
It is straightforward to verify that $\mathcal{T}_{u^{\ast}}=[-1,1]$. From this it follows that $A_{+}(u^{\ast})=\varnothing=A_{-}(u^{\ast})$, and finally that $\mathcal{T}^{nb}_{u^{\ast}}=[-1,1]$.
The case of $X^{\ast}=\mathbf{S}^{1}$ follows similarly and directly applying [16, Lemma 3.4.2]. ∎
We state the following straightforward lemma without proof.
@@
Let $u\colon X\to\mathbb{R}$ be the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. The sets $\Gamma_{u}$, $\Gamma_{u}^{-}$, $\mathcal{T}_{u}$, $A_{+}(u)$, $A_{-}(u)$, $\Gamma_{u}(x)$, and $\Gamma_{u}^{-}(x)$ are $G$-invariant. That is if a point is contained in any of these sets, then their orbit is contained in the set.
We now obtain the following corollary from Lemma 3.3.
Corollary 3.5.
Let $G$ be a Lie group acting by isometries on a metric space $(X,d)$ so that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, or $\mathbf{S}^{1}$. Let $u\colon X\to\mathbb{R}$ the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$. Then the transport set $\mathcal{T}^{nb}_{u}$ is open.
Proof.
By Lemma 3.3 we have that $(\mathcal{T}^{nb}_{u})^{\ast}$ is $[-1,1]$, $[0,\infty)$ or $\mathbb{R}$, if $X^{\ast}$ is isometric to $[-1,1]$ $[0,\infty)$ or $\mathbb{R}$ respectively, and $(\mathcal{T}^{nb}_{u})^{\ast}$ is $\mathbf{S}^{1}\setminus\{-x_{0}^{\ast}\}$ if $X^{\ast}$ is isometric to $\mathbf{S}^{1}$. Moreover, $(\mathcal{T}^{nb}_{u})^{\ast}=\mathcal{T}^{nb}_{u^{\ast}}$.
In all cases, $y\in\mathcal{T}^{nb}_{u}$ we can find $\varepsilon>0$ small enough so that $B_{\varepsilon}(y^{\ast})\subset\mathcal{T}^{nb}_{u^{\ast}}$. We observe that for every $x^{\ast}\in B_{\varepsilon}(y^{\ast})$ there exists $x\in\pi^{-1}(B_{\varepsilon}(y^{\ast}))$ with $x\in\mathcal{T}^{nb}_{u}$. Thus we we have $\pi^{-1}(x^{\ast})=G(x)\subset\mathcal{T}^{nb}_{u}$, and $G(x)=\pi^{-1}(x^{\ast})\subset\pi^{-1}(B_{\varepsilon}(y^{\ast}))$. This implies that $\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset\mathcal{T}^{nb}_{u}$. But since we have that the topology of the metric $d^{\ast}$ agrees with the quotient topology, we have that $\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset X$ is open. This implies that there exists $\delta>0$ such that $B_{\delta}(y)\subset\pi^{-1}(B_{\varepsilon}(y^{\ast}))\subset\mathcal{T}^{nb}_{u}$. Thus we conclude that $\mathcal{T}^{nb}_{u}$ is open. ∎
From this we easily conclude the following corollary.
Corollary 3.6.
Let $G$ be a compact Lie group acting by isometries on a metric space $(X,d)$ so that $X^{\ast}$ is isometric to $[-1,1]$, $[0,\infty)$, or $\mathbb{R}$. For $u\colon X\to\mathbb{R}$ the $G$-invariant Lipschitz function given by $u(y)=d(G(x_{0}),y)$, the transport set $\mathcal{T}^{nb}_{u}$ is $X$.
With these results established we are in position to prove Theorem C.
Proof of Theorem C.
Without loss of generality, we may assume that $X^{\ast}=[-1,1]$. Fix $x_{0}\in X$ and take $\delta>0$, such that $B_{\delta}(x_{0})\subset\mathcal{T}^{nb}_{u}$, and $B_{\delta}(G(x_{0}))\setminus G(x_{0})$ consists of only principal orbits, that is, $\pi(B_{\delta}(G(x_{0}))\setminus G(x_{0}))$ is homeomorphic to $(0,\delta)$ or $(-\delta,0)\cup(0,\delta)$. We claim that
(Slice) $S_{x_{0}}:=(\Gamma^{-}_{u}(x_{0})\cap B_{\delta}(x_{0}))\cup\{x_{0}\}.$
is a slice through $x_{0}$. In order to prove so, we check each item in the definition of Slice as recalled in Section 2.
Point (i). Take $g\in G_{x_{0}}$, and $y\in S_{x_{0}}$. Since $u(x_{0})=0$ we conclude that $(x_{0},y)\in\Gamma_{u}^{-}$ and in turn that $(y,x_{0})\in\Gamma_{u}$. That is, $d(G(x_{0}),y)=u(y)=d(x_{0},y)$. This implies that
$d(x_{0},gy)=d(g^{-1}x_{0},y)=d(x_{0},y)=d(G(x_{0}),y)=d(G(x_{0}),gy).$
Therefore, $(gy,x_{0})\in\Gamma_{u}$ which gives $(x_{0},gy)\in\Gamma^{-}_{u}$, so that $gy\in\Gamma^{-}_{u}(x_{0})$. Moreover, $d(x_{0},gy)=d(x_{0},y)<\delta$. Finally, observe that all the sets $\mathcal{T}_{u}$, $A_{+}$ and $A_{-}$ are $G$-invariant. Thus $\mathcal{T}^{nb}_{u}$ is $G$-invariant, and therefore, $gy\in S_{x_{0}}$.
Point (ii). We begin by observing that under $R_{u}$, the equivalence class of $y\in S_{x_{0}}$ is the unique geodesic ray $\gamma_{y}$ starting at $x_{0}$ realizing the distance $d(G(x_{0}),y)=d(x_{0},y)$, i.e. realizing the distance from $y$ to $G(x_{0})$.
Now assume that for some $g\in G$ we have $gS_{x_{0}}\cap S_{x_{0}}\neq\emptyset$. That is, there exist $x,y\in S_{x_{0}}$ such that $gx=y$. Then we let $\gamma_{x}$ and $\gamma_{y}$ be the unique geodesic segments starting at $x_{0}$, with end points $x$ and $y$, and realizing the distances $d(G(x_{0}),x)=d(x_{0},x)$ and $d(G(x_{0}),y)=d(x_{0},y)$ respectively. Observe that $g\gamma_{x}$ is also a geodesic segment starting at $gx_{0}$ with endpoint $gx=y$ and realizing the distance $d(gx_{0},y)=d(gx_{0},gx)=d(G(x_{0}),gx)=d(G(x_{0}),y)=d(x_{0},y)$. Thus, as the equivalence class of $y$ in $\mathcal{T}^{nb}_{u}$ under $R_{u}$ is the unique geodesic joining $y$ to $G(x_{0})$, realizing the distance $d(G(x_{0}),y)$, this implies that $g\gamma_{x}=\gamma_{y}$ and in particular has to also start at $x_{0}$. Whence $gx_{0}=x_{0}$, and thus $g\in G_{x_{0}}$.
Point (iii). Observe that we have a principal bundle $\rho\colon G\to G/G_{x_{0}}$, where we consider the right action of $G_{x_{0}}$ on $G$. Set $H=G_{x_{0}}$. This implies the existence of an open neighborhood $U\subset G/H$ containing $eH$, such that $\rho^{-1}(U)$ is equivarantly homeomorphic to $H\times U$. This further implies the existence of a continuous cross-section $\chi\colon U\to G$. We now define a map $F\colon U\times S_{x_{0}}\to X$ by $F(gH,s)=\chi(gH)s$. We claim that $F$ is a homeomorphism onto its image.
To prove the injectivity of $F$, consider $(g_{1}H,s_{1})$, $(g_{2}H,s_{2})\in U\times S_{x_{0}}$, such that $\chi(g_{1}H)s_{1}=\chi(g_{2}H)s_{2}$. This implies $\chi(g_{2}H)^{-1}\chi(g_{1}H)s_{1}=s_{2}\in S_{x_{0}}$. By Point (ii), this implies that $\chi(g_{2}H)^{-1}\chi(g_{1}H)\in H$, and thus for some $g\in H$, we have $g\chi(g_{2}H)=\chi(g_{1}H)$. Thus $g_{2}H=\rho(g\chi(g_{2}H))=\rho(\chi(g_{1}H))=g_{1}H$. This last part implies that $s_{1}=s_{2}$. Therefore $F$ is injective.
We now show that $F$ is onto and open. Take $V\subset U$ open, and consider the homeomorphism $\tilde{\alpha}_{x_{0}}\colon G/H\to G(x_{0})$ guaranteed by Theorem 2.26.
Consider $y\in B_{\delta}(\tilde{\alpha}_{x_{0}}(V))\subset X$. Then there exists an $gH\in V$ such that for $gx_{0}\in G(x_{0})$ we get $\delta>d(G(x_{0}),y)=d(gx_{0},y)$. Observe that we have $\chi(gH)x_{0}=\alpha_{x_{0}}(\chi(gH))=\tilde{\alpha}_{x_{0}}(\rho(\chi(gH)))=\tilde{\alpha}_{x_{0}}(gH)=gx_{0}$. Denote by $\gamma_{y}$ the minimizing geodesic realizing the distance between $gx_{0}$ and $y$. Then the geodesic $\chi(gH)^{-1}\gamma_{y}$ joining $\chi(gH)^{-1}y$ to $x_{0}$ has length $d(\chi(gH)^{-1}y,x_{0})=d(y,gx_{0})=d(G(x_{0}),y)=d(G(x_{0}),\chi(gH)^{-1}y)$. Thus $\chi(gH)^{-1}y$ is a point in $\Gamma_{u}(x_{0})$ and in $B_{\delta}(x_{0})$. Therefore $\chi(gH)^{-1}y$ is a point in $S_{x_{0}}$. Thus we have that $(gH,\chi(gH)^{-1}y)\in V\times S_{x_{0}}$ with $F(gH,\chi(gH)^{-1}y)=y$. This implies that $B_{\delta}(\tilde{\alpha}_{x_{0}}(V))\subset F(U\times S_{x_{0}})$. That is $F$ is an open bijective continuous map onto its image, and thus it is an homeomorphism onto its image. ∎
Remark 3.7.
We point out that in the previous proofs we only need to be able to define a metric $d^{\ast}$ on $X^{\ast}$, and that $(X^{\ast},d^{\ast})$ is isometric to $\mathbb{R}$, a closed ray, a closed interval, or $\mathbf{S}^{1}$ with the length metric induced from $\mathbb{R}^{2}$. When $G$ is non-compact, by asking that the action be proper, i.e. the map $\bar{A}\colon G\times X\to X\times X$ is a proper map, is a sufficient condition so that there exists a $G$-invariant metric $d$ on $X$ such that $d^{\ast}$ is well defined. This follows from the fact that the isotropy groups of proper actions are compact, that the orbits are closed, and the orbit space is Hausdorff (the orbits cannot get arbitrarily close to each other).
One can drop the assumption on the dimension of the orbit space in Corollary 3.5 and Theorem C to obtain the same conclusions if the orbit space is isometric to an Alexandrov space, as we now show. Note however that we are still not assuming any curvature bound on $X$.
Theorem 3.8.
Let $G$ act by isometries on $(X,d)$ a metric space, and assume that $(X^{\ast},d^{\ast})$ is isometric to an Alexandrov space. Moreover, assume that there exists $\mathcal{M}\subset X^{\ast}$ an open subset such that for $x_{0}^{\ast}\in\mathcal{M}$ the distance of the cut locus $\mathcal{C}(x_{0}^{\ast})$ to $x_{0}^{\ast}$ is positive. Then the set $\mathcal{T}^{nb}_{u^{\ast}}$ is open, where $u^{\ast}\colon X^{\ast}\to\mathbb{R}$ is defined as $u^{\ast}(x^{\ast})=d(x_{0}^{\ast},x^{\ast})$. Moreover, for any point $x\in\pi^{-1}(x_{0})$ we have a slice.
Proof.
We fix $x_{0}^{\ast}\in\mathcal{M}$, and consider $u^{\ast}\colon X^{\ast}\to\mathbb{R}$ given by $u^{\ast}(y^{\ast})=d^{\ast}(x_{0}^{\ast},y^{\ast})$. We observe that given $y^{\ast}\in A_{\pm}(u^{\ast})$, there exists two distinct geodesics $\gamma_{1}^{\ast}$ and $\gamma_{2}^{\ast}$ of $(X^{\ast},d^{\ast})$ starting at $x_{0}^{\ast}$, passing through $y^{\ast}$, such that there exist points $w^{\ast}\in\gamma_{1}^{\ast}$ and $v^{\ast}\in\gamma_{2}^{\ast}$ for which the shortest geodesic $\gamma_{3}^{\ast}$ between them does not contain $x_{0}^{\ast}$. This implies that $y^{\ast}$ is a critical point of $u^{\ast}$ by [79, p. 378]. Then our hypothesis guarantee that there is a sufficiently small ball $B$ centered at $x_{0}^{\ast}$ such that $B\cap A_{\pm}(u^{\ast})=\varnothing$. From this our conclusion follows.
The existence of the slice is proven in an analogous way to the proof of Theorem C by considering $\delta>0$ small enough so that $\pi(B_{\delta}(G(x_{0})))\subset B^{\ast}$. ∎
Remark 3.9.
The condition asking $x_{0}^{\ast}$ to be isolated from its cut locus $\mathcal{C}(x_{0}^{\ast})$ is a not necessary for the existence of the slice. For example if $D\subset\mathbb{R}^{2}$ is a convex domain, then we consider $X$ to be its double which is an Alexandrov space with curvature $\geq 0$. For $x\in\partial D\subset X$, we have that the cut locus is $\mathcal{C}(x)=\partial D\setminus\{x\}$ (see [71, Section 1, Example 2]). Thus $x$ is not isolated from its cut locus. Considering $G$ any compact Lie group with a left-invariant metric, the space $Y:=G\times X$ with the product metric is an Alexandrov space, whose orbit space is $X$, and for any point $(g,x)\in G\times X$, $S_{(g,x)}:=\{g\}\times X$ is a slice.
3.1. Kleiner’s Lemma
Here we show that a result analogous to Kleiner’s Lemma for Riemannian manifolds [54, Lemma 2.2.2] and Alexandrov spaces [26, Lemma 2.1] holds for cohomogeneity one geodesic spaces:
Lemma 3.10 (Kleiner’s Lemma).
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, $\mathbf{S}^{1}$. Then for a minimal geodesic $\gamma\colon[0,d]\to X$ such that $d(\gamma(0),\gamma(d))=d^{\ast}(\gamma(0)^{\ast},\gamma(d)^{\ast})$, we have that for $t\in(0,d)$ the isotropy group $G_{\gamma(t)}$ is a subgroup of $G_{\gamma(0)}$ and $G_{\gamma(d)}$.
Proof.
We point out that for any $t\in[0,d]$ there exists a subinterval $(t-\varepsilon,t+\varepsilon)\subset[0,d]$ such that the geodesic $\gamma|_{(t-\varepsilon,t+\varepsilon)}$ is contained in the slice $S_{\gamma(t)}$ given by Theorem C. Thus we can apply an analogous argument to the one given in the proof of [26, Lemma 2.1]:
Without loss of generality, assume that there exists $t_{0}\in(0,d)$ such that there exits $g\in G_{\gamma(t_{0})}$ with $g\not\in G_{\gamma(d)}$. The case of assuming that $g\not\in G_{\gamma(0)}$ is analogous. Then, there exists possibly a new $t_{0}^{\prime}$ and a subinterval $[t_{0}^{\prime}-\delta,t_{0}^{\prime}+\delta]$ such that the geodesic $\gamma|_{[t_{0}^{\prime}-\delta,t_{0}^{\prime}+\delta]}$ is in the slice $S_{\gamma(t_{0}^{\prime})}$ and $g\gamma(t_{0}^{\prime}+\delta)\neq\gamma(t_{0}^{\prime}+\delta)$. But this gives a branching geodesic contained in the slice $S_{\gamma(t_{0}^{\prime})}$ which is a contradiction. ∎
With this we prove that the isotropy group is constant in the interior of the orbit space:
Lemma 3.11 (Principal Isotropy Lemma).
Let $(X,d)$ be a geodesic space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to $[-1,1]$, $[0,\infty)$, $\mathbb{R}$, $\mathbf{S}^{1}$. All the orbits corresponding to the interior of $X^{\ast}$ have constant isotropy subgroup.
Proof.
Consider $x\in X$, such that $x^{\ast}\in\mathrm{int}(X^{\ast})$. Then there exists $y^{\ast}\in\mathrm{int}(X^{\ast})$, such that $x$ is in the slice $S_{y}$ for some $y\in\pi^{-1}(y^{\ast})$ for $\pi\colon X\to X^{\ast}$. In particular $x$ is in the interior of a minimizing geodesic starting at $y$ realizing the distance between the orbits. Then by Lemma 3.10, it follows that $G_{x}\subset G_{y}$. But we can apply the same idea to $y$, now taking the slice through $x$. Thus we get that $G_{y}=G_{x}$, and this is true for all points in $\pi^{-1}(\mathrm{int}(X^{\ast})$. ∎
3.2. Group Diagram
We now show a relation between the isotropy subgroups for an action of a compact Lie group $G$ on a geodesic space $(X,d)$ by isometries. From this relation we obtain a group diagram for actions of cohomogeneity one by isometries on geodesic spaces.
Lemma 3.12.

### v3

@@ 3. Slice Theorem for cohomogeneity one geodesic spaces
    Let $y_{n},n\in\mathbb{N}$ be a converging sequence of points in $S_{x_{0}}$ such that the limit $y\in G(S_{x_{0}})$. As the function $d(x_{0},\cdot)\colon X\rightarrow[0,\infty)$ is continuous we have that $d(x_{0},y_{n})\rightarrow d(x_{0},y)$ as $n\rightarrow\infty$. Also, notice that $d(x_{0},y_{n})=d(G(x_{0}),y_{n})=d^{\ast}(x_{0}^{\ast},y_{n}^{\ast})\rightarrow d^{\ast}(x_{0}^{\ast},y^{\ast})$ as $n\rightarrow\infty$.
    Therefore $d(x_{0},y)=d^{\ast}(x_{0}^{\ast},y^{\ast})=d(G(x_{0}),y)$.
@@
    Just notice that
    $G(S_{x_{0}})=\bigcup_{gx_{0}\in x_{0}^{\ast}}S_{gx_{0}}=G(B_{\delta}(gx_{0}))$
    with the set in the last equality clearly being an open neighborhood of the orbit $x_{0}^{\ast}$.
@@
    Take $g\in G_{x_{0}}$ and $y\in S_{x_{0}}$. Then $d(x_{0},y)=d(x_{0},gy)$. Take $hx_{0}\in\pi^{-1}(x_{0}^{\ast})$ such that $d(hx_{0},gy)=d(G(x_{0}),gy)$ and observe that
    $d(G(x_{0}),y)=d(G(x_{0}),gy)=d(g^{-1}hx_{0},y).$
    Since $y^{\ast}\in intX^{\ast}$ by ?THM? LABEL:proposition.uniquerealiser this implies that $g^{-1}hx_{0}=x_{0}$ so then $h\in G_{x_{0}}$ and $gy\in S_{x_{0}}$.
@@
    First, observe that for any $g\in G$ $gS_{x_{0}}=S_{gx_{0}}$. Now, take a point $y\in S_{gx_{0}}\cap S_{x_{0}}$ and so
    $d(gx_{0},y)=d(x_{0},y)=d(G(x_{0}),y).$
    Again, by ?THM? LABEL:proposition.uniquerealiser it follows that $gx_{0}=x_{0}$ which means that $g\in G_{x_{0}}$.
As all of the conditions are satisfied, we conclude that $S_{x_{0}}$ is indeed a slice at $x_{0}$. ∎
Remark 3.15.
Due to ?THM? LABEL:proposition.uniquerealiser, we have that $S_{x_{0}}$ is a union of horizontal geodesics starting which have $x_{0}$ as the single common point. Thus $S_{x_{0}}$ is homeomorphic to a cone.
If we allow the dimension of the quotient space $X^{\ast}$ to be larger than one, we can still get a slice but we require stronger assumptions on both $X$ and $X^{\ast}$.
Theorem 3.16.
Let $G$ act by isometries on a non-branching space $(X,d,\mathfrak{m})$ and assume that $(X^{\ast},d^{\ast})$ is an Alexandrov space. Moreover, assume that there exists $\mathcal{M}\subseteq X^{\ast}$ an open set such that for $x_{0}^{\ast}\in\mathcal{M}$ the distance of the cut locus $\mathcal{C}(x_{0}^{\ast})$ to $x_{0}^{\ast}$ is positive. Then, for every point $x\in x_{0}^{\ast}$ we have a slice.
@@
Let $x_{0}^{\ast}$ and take $0<\delta<d^{\ast}(x_{0}^{\ast},\mathcal{C}(x_{0}^{\ast}))$. For an orbit $y^{\ast}\in B_{\delta}(x_{0}^{\ast})$ take some $y\in\pi^{-1}(y^{\ast})$. Now, there exists some $gx_{0}\in\pi^{-1}(x_{0}^{\ast})$ such that $d(gx_{0},y)=d(G(x_{0}),y)$.
Noting that our space is non-branching and that we can find a geodesic $\gamma\in\operatorname{Geo}(X)$ such that $\gamma_{0}=gx_{0}$ and $\gamma_{t}=y$ for some $t\in(0,1)$ we can conclude that
$\arg\min\{d(gx_{0},y)\,|\,gx_{0}\in\pi^{-1}(x_{0}^{\ast})\}=\{gx_{0}\}.$
Furthermore, by arguing as in ?THM? LABEL:lemma.projection.horizontalgeo we can check that $\gamma$ is a horizontal geodesic. Once we have this we can define
$S_{x_{0}}=\{y\in X\,|\,d(x_{0},y)=d(G(x_{0}),y)\}\cap B_{\delta}(x_{0})$
and proceed as in the proof of ?THM? LABEL:MT:_Slice_Theorem_Cohomogenity_one to conclude that it is indeed a slice. ∎
Remark 3.17.
In the previous Theorem we needed to impose further restrictions on the geodesics of $X$. This is because if the dimension of the quotient space $X^{\ast}$ is larger than or equal to $2$, it could happen that ?THM? LABEL:proposition.nonbranchinghorizontalgeodesic does not hold. This is because some horizontal branching geodesic $\gamma$ could very well be undetected as there is no guarantee that taking $\{g\gamma\mid g\in G\}$ would yield something of positive measure.
Remark 3.18.
The condition asking $x_{0}^{\ast}$ to be isolated from its cut locus $\mathcal{C}(x_{0}^{\ast})$ is a not necessary for the existence of the slice. For example if $D\subset\mathbb{R}^{2}$ is a convex domain, then we consider $X$ to be its double which is an Alexandrov space with curvature $\geq 0$. For $x\in\partial D\subset X$, we have that the cut locus is $\mathcal{C}(x)=\partial D\setminus\{x\}$ (see [75, Section 1, Example 2]). Thus $x$ is not isolated from its cut locus. Considering $G$ any compact Lie group with a left-invariant metric, the space $Y:=G\times X$ with the product metric is an Alexandrov space, whose orbit space is $X$, and for any point $(g,x)\in G\times X$, $S_{(g,x)}:=\{g\}\times X$ is a slice.
3.1. Cohomogeneity one Group diagrams
We now show a relation between the isotropy subgroups for an action of a compact Lie group $G$ on a geodesic space $(X,d)$ by isometries. From this relation we obtain a group diagram for actions of cohomogeneity one by isometries on essentially non-branching spaces. Given an action of the compact Lie group $G$ on a metric space $(X,d)$ such that $(X^{\ast},d^{\ast})$ is homoemorphic to [1,1][-1,1], a group diagram is a tuple $(G,K_{-},K_{+},H)$ where $H,K_{-},K_{+}<G$ such that the following diagram of inclusions holds:
[…]
Theorem 3.19.
Let $(X,d,\mathfrak{m})$ be an essentially non-branching space, and $G$ a compact Lie group acting by isometries, such that $(X^{\ast},d^{\ast})$ is isometric to [1,1][-1,1]. Then we have three isotropy subgroups $H$, $K_{-}$, and $K_{+}$, whose isotropy types correspond to the orbits in the interior (1,1)(-1,1), and the end points $\{-1\}$ and $\{+1\}$ respectively. Moreover, $H$ is a subgroup of both $K_{-}$ and $K_{+}$.
Proof.
We fix $\gamma\colon[0,1]\to X$ a geodesic realizing the distance between the orbits, guaranteed to exist by ?THM? LABEL:proposition.nonbranchinghorizontalgeodesic. Then by Kleiner’s ?THM? LABEL:L:_Kleiners_Lemma the conclusions follow, by taking $H:=G_{\gamma_{0}}$, and $K_{\pm}=G_{\gamma_{\pm 1}}$. ∎
The tuple $(G,H,K_{+},K_{-})$ obtained in ?THM? LABEL:T:_Principal_isotropy_is_subgroup_of_singular_isotropy depends a priori on a choice of the geodesic. In the following lemma we show how two tuples associated to two different geodesics relate to each other.